OpenStudy (anonymous):

How do I find the vertex of the parabola y=x(squared) + 8x +5

OpenStudy (anonymous):

\(-\frac{b}{2a}\) is the first coordinate of the vertex of \(ax^2+bx+c\) in your case \(a=1,b=8\)

OpenStudy (jdoe0001):

\(\bf \text{vertex of a parabola} \qquad \qquad \left(-\cfrac{b}{2a}, c-\cfrac{b^2}{4a}\right)\)